The lifetime T in hours of a system (with some redundancy) has pdf fr(t) Cte 110)(t). What is the probability that the system will last for at least 5 hours?
Answer #1

here for above to be valid: \int_{0}^{\infty } f(t) dt =1

\int_{0}^{\infty } f(t) dt =\int_{0}^{\infty } Cte-t/2 dt =-C*(2te-t/2-4e-t/2)|\infty0 =4C =1

therefore C =1/4

hence P(system last for at least 5 hours) =P(T>5)

=\int_{5}^{\infty } f(t) dt =\int_{5}^{\infty } (1/4)te-t/2 dt =-(1/4)*(2te-t/2-4e-t/2)|\infty5 =0.2873